OpenStudy (logo):

Suppose a parabola has an axis of symmetry at x = -2, a minimum height at -6, and passes through the point (0, 10). Write the equation of the parabola in vertex form.

OpenStudy (breanna1999):

a(x-h)^2 + k is the vertex form

OpenStudy (breanna1999):

a(x+2)^2+k is the next step

OpenStudy (dumbcow):

vertex = (-2,-6) = (h,k) plug in point (0,10) for (x,y) solve for a

OpenStudy (logo):

how is it a(x+2) and not a(x-2) ??

OpenStudy (dumbcow):

x-h , h = -2, x-(-2) = x+2

OpenStudy (logo):

so 10 = a(0 + 2) ^2 - 6 10 = a(2)^2 - 6 10 = a * 4 - 6 10 = a* -2 a = -2 -10 a = 12 y = 12( x + 2)^2 - 6

OpenStudy (logo):

@dumbcow

OpenStudy (logo):

I mean a = -12 not 12

OpenStudy (dumbcow):

almost, go back to step: 4a -6 = 10 you cant add the 4 and -6 add 6 to both sides 4a = 16

OpenStudy (logo):

oh ok so then a = 4 y = 4(x+2)^2 - 6? @dumbcow